A train travels 160km in the same time that a plane covers 448km. If the speed of the plane is 10km per hr less than3 times the speed of the train, find both speeds.
Answer by stanbon(75887)

A train travels 160km in the same time that a plane covers 448km. If the speed of the plane is 10km per hr less than3 times the speed of the train, find both speeds.
---------------
Train DATA:
distance = 160 km ; rate = x km/h ; time = 160/x hrs
--------------
Plane DATA:
distance = 448 km ; rate = 3x-10 km/h; time = 448/(3x-10) hrs
--------------
Equation:
time = time
160/x = 448/(3x-10)
160(3x-10) = 448x
480x - 1600 = 448x
32x = 1600
x = 50 km/h (train rate)
3x-10 = 140 km/h (plane rate)
==================
Cheers,
Stan H.